OpenStudy (anonymous):

Let \( \pi(n)\) be the number of primes less or equal to n. Show that \[ n^{\pi(2n)-\pi(n)}<4^{n} \]

OpenStudy (anonymous):

What can we use as the basis for our proof? May we assume PNT?

OpenStudy (anonymous):

You can if you want. There is an elementary proof of it not involving the PNT

OpenStudy (anonymous):

The PNT is for large enough n. Here one has to prove the inequality fo each n.

OpenStudy (anonymous):

If there's an elementary proof, I have no idea how to start this... will think over it, for tomorrow... seems pretty interesting.

OpenStudy (anonymous):

*

OpenStudy (anonymous):

Hint \[ (1+1)^{2n}=4^n \]

OpenStudy (anonymous):

\[ 4^n=(1+1)^{2n}> {2n \choose n} \]

OpenStudy (anonymous):

oh very nice for \(n\le p_k\le 2n\) \[\prod p_k | \frac{(n+1)(n+2)...(2n)}{n!}={2n \choose n} \Rightarrow\prod p_k <{2n \choose n}\]\[n^{\pi(2n)-\pi(n)}< \prod p_k<\binom{2n}{n}< 2^{2n}=4^{n} \]

OpenStudy (anonymous):

Bah, humbug, you got me... but you were right, there *was* a nice proof! Figured it out sometime earlier during class.